# A Panchromatic View of Late-time Shock Power in the Type II Supernova 2023ixf

### v2

@@ Abstract
We present multi-wavelength observations of the type II supernova (SN II) 2023ixf during its first two years of evolution. We combine ground-based optical/NIR spectroscopy with Hubble Space Telescope (HST) far- and near-ultraviolet spectroscopy and James Webb Space Telescope (JWST) near- and mid-infrared photometry and spectroscopy to create spectral energy distributions of SN 2023ixf at +374 and +620 days post-explosion, covering a wavelength range of 0.1-30 m. The multi-band light curve of SN 2023ixf follows a standard radioactive decay decline rate after the plateau until 500 days, at which point shock powered emission from ongoing interaction between the SN ejecta and circumstellar material (CSM) begins to dominate. This evolution is temporally consistent with 0.3-10 keV X-ray detections of SN 2023ixf and broad “boxy” spectral line emission from reprocessing of shock luminosity in a cold dense shell located between forward and reverse shocks. Using the expected absorbed radioactive decay power and the detected X-ray luminosity, we quantify the total shock powered emission at the +374 and +620 day epochs and find that it can be explained by nearly complete thermalization of the reverse shock luminosity as SN 2023ixf interacts with a continuous, ”wind-like” CSM with a progenitor mass-loss rate of M⊙ yr-1 ( km/s). Additionally, we construct multi-epoch spectral models from the non-LTE radiative transfer code CMFGEN, which contain radioactive decay and shock powers, as well as dust absorption, scattering, and emission. We find that models with shock powers of erg s-1 and of silicate dust in the cold dense shell and/or inner SN ejecta can effectively reproduce the global properties of the late-time (¿300 days) UV-to-IR spectra of SN 2023ixf.
[…]
@@ 1 Introduction
Supernova (SN) 2023ixf is a type II supernova (SN II) that occurred on 2023-05-18 in Messier 101 at a distance of Mpc (Riess et al., 2022). SN 2023ixf is the closest SN of the last decade and, consequently, was reported to the Transient Name Server (TNS) within 1 day of first light by Koichi Itagaki (Itagaki, 2023) on 2023 May 19 17:27:15 (MJD 60083.73). The earliest reported detection of SN 2023ixf was on MJD 60082.85 at mag (-12.1 mag) in -band following a deep upper limit of mag ( mag) on MJD 60082.66 (Mao et al., 2023; Li et al., 2024). A conservative estimate of the time of first light is MJD , which is adopted throughout this work.
The early-time spectra of SN2023ixf exhibit emission lines from H i, He i/ii, C iv, and N iii/iv/v, superimposed on a hot, blue spectral continuum (Jacobson-Galán et al., 2023; Bostroem et al., 2023; Teja et al., 2023; Zhang et al., 2023). These lines, commonly referred to as “flash” or “IIn-like” features, originate from the sustained photoionization of dense circumstellar material (CSM) located ahead of the forward shock (Gal-Yam et al., 2014; Khazov et al., 2016; Yaron et al., 2017; Dessart et al., 2017; Terreran et al., 2022; Bruch et al., 2021, 2023; Dessart & Jacobson-Galán, 2023; Jacobson-Galán et al., 2024a). Much like Type IIn supernovae, the emission line profiles consist of a narrow core accompanied by broad Lorentzian “wings” that extend up to 1000 , a signature of electron scattering within the ionized, optically thick CSM (Chugai, 2001; Dessart et al., 2009, 2017; Huang & Chevalier, 2018). Additionally, SN 2023ixf has the earliest and highest cadence UV spectroscopy (Teja et al., 2023; Zimmerman et al., 2024), spectropolarimetry (Vasylyev et al., 2023; Singh et al., 2024; Shrestha et al., 2025; Vasylyev et al., 2025) and high resolution spectroscopy (Smith et al., 2023; Zimmerman et al., 2024; Dickinson et al., 2025) of any CSM-interacting SN II to date.
@@ 2.1 Photometric Observations
In addition to our observations, we include -band photometry from the Zwicky Transient Facility (ZTF; Bellm et al. 2019; Graham et al. 2019) forced-photometry service (Masci et al., 2019). We also include multi-band photometry presented in Jacobson-Galán et al. (2023); Van Dyk et al. (2024a); Hsu et al. (2024); Zimmerman et al. (2024); Singh et al. (2024) in our analysis. Furthermore, we include JWST MIRI imaging of SN 2023ixf from 2024-03-15 (JWST-GO-3921, PI: Fox) and 2025-01-27 (JWST-GO-5290, PI: Ashall), which was reduced using spacephot photometry package.333https://space-phot.readthedocs.io/en/latest/ The complete UVOIR light curve of SN 2023ixf is shown in Figure 2.
@@ 2.2 Spectroscopic Observations
Far- and near-UV spectra of SN 2023ixf were obtained on 2024-03-24 ( days) and 2025-01-26 ( days) with the Hubble Space Telescope (HST) using the Multi-Anode MicroChannel Array (MAMA) detectors on the Space Telescope Imaging Spectrograph (STIS) through HST programs 17497 (PI Valenti) and 17772 (PI Bostroem). Standard bias subtraction, flat-fielding and cosmic ray rejection were performed automatically before downloading from the Mikulski Archive for Space Telescopes (MAST). Wavelength calibration, flux calibration and 1-D extraction was performed using the calstis pipeline routine. Near- and mid-infrared spectroscopy of SN 2023ixf were obtained on 2024-01-26, 2024-05-26, and 2025-01-08 ( days) (JWST-DD-4575, JWST-GO-5290; PI: Ashall) with the James Webb Space Telescope (JWST) using NIRSpec with G235M-F170LP and G395M-F290LP grating/filter combinations (Jakobsen et al., 2022) and MIRI/LRS (Kendrew et al., 2015). These data were reduced using the jwst555https://github.com/spacetelescope/jwst pipeline, which performs standard bias subtraction, flat-fielding, wavelength and flux calibrations, and spectral extraction. The complete optical/NIR spectral sequence is shown in Figure 3 and a log of all spectral observations utilized in this paper is provided in Table A1. All photometric and spectroscopic data will be made public online.666https://github.com/wynnjacobson-galan/SN2023ixf
@@ 3.1 Radioactive Decay Power
After a 83 day light curve photospheric phase (see Figure 2), the photometric evolution of SN 2023ixf became dominated by the radioactive decay from the standard Fe chain (Arnett, 1982). Modeling of the nebular decline rate has suggested and incomplete -ray trapping with a characteristic timescale of days (Zimmerman et al., 2024; Singh et al., 2024; Li et al., 2025). To confirm these parameters, we construct two variations of the post-plateau bolometric light curve, which are then fit with analytic formalisms for radioactive decay power. First, we construct a pseudobolometric light curve using filters and fit this light curve with the pseudobolometric light curve of SN 1987A derived with the same filter combination (e.g., see Eqn. 1 of Jacobson-Galán et al. 2025), which yields and days. We note that consistent values are found when this method is applied to the more confined filter combination. Second, we construct a pseudobolometric light curve from the combination of filters and which covers m. Given the emergence of shock powered emission at late-times (§3.2), we model this pseudobolometric light curve at days and find and days. We chose to exclude UV and mid-IR photometry in this light curve given that (1) of SN II luminosity is expected to be emitted between m at days (Dessart, 2025a) if it arises from decay power only and (2) the UV excess from persistent shock powered emission shown in Figure 5 would lead to an overestimate of . Based on the average of both modeling attempts, we choose to adopt and days when calculating the contribution of absorbed radioactive decay power to the observed luminosity of SN 2023ixf.
@@ 3.2 Shock Powered Emission
As shown in Figures 2 & 5, an excess emission beyond what is expected from pure radioactive decay first becomes apparent at UV wavelengths as the Swift-UVOT light curve becomes flat at days and remains bright at mag above the host galaxy background – these observations are also consistent with observed emission in the near- and far-UV at days (Bostroem et al., in prep). Host galaxy subtracted optical and NIR photometry shows a similar flattening at days and the pseudobolometric light curve (e.g., Fig. 5) deviates from a radioactive decay decline rate as shock powered emission from persistent SN ejecta-CSM interaction becomes the dominant power source in SN 2023ixf. Here we define “shock powered emission” as the emergent UVOIR luminosity from reprocessed X-ray luminosity from the radiative RS i.e., all additional UVOIR luminosity that does not come from absorbed radioactive decay power. The presence of shock powered emission in SN 2023ixf is also confirmed by the spectral evolution shown in Figure 3 as the forbidden emission lines observed at year fade and prominent boxy H, H, P, He i and [Fe ii] emission become dominant at days – this transition in H is shown in Figure 6. The presence of these emission line profiles is direct evidence for a radiative reverse shock that injects power into both the CDS and the inner ejecta (Nymark et al., 2006; Chevalier & Fransson, 2017; Dessart & Hillier, 2022). Similarly, prominent Ly and Mg ii emission from the CDS are observed at UV wavelengths (e.g., Fransson et al. 2005; Dessart et al. 2023, Bostroem et al. in prep). Furthermore, evidence for ongoing CSM interaction is given by persistent X-ray and radio emission, both of which imply a mass loss rate of M⊙ yr-1 at shock radii cm (Panjkov et al., 2024; Chandra et al., 2024; Nayana et al., 2025). We note that the multi-wavelength evolution of SN 2023ixf confirms that shock powered emission has always been present as the SN ejecta continues to interact with the likely steady-state wind ahead of the forward shock. However, it is overall unavoidable to have some radiation produced via dissipation of kinetic energy through shocks, and this would be true in SN 2023ixf, as in all SNe, even in the absence of X-ray and radio detections. Nevertheless, this shock powered emission source has remained subdominant compared to radioactive decay power until days in optical/IR bands and its emergence at these late times is not the result of SN ejecta interaction with some detached shell of CSM.
@@
where is the ejecta velocity at the RS in units of and is the CSM density power-law index. For a radiative cooling RS and a density profile, the time variable above disappears and the luminosity decreases as , which then goes as for and given the time dependence of the ejecta velocity (e.g., see Eqn. 2.2 in Fransson et al. 1996). We adopt M⊙ yr-1() and choose an ejecta velocity of , which covers the maximum velocity observed in the “boxy” H profile (e.g., Fig. 6). Applying Eqn. 3.14 in Fransson et al. (1996), we confirm that the RS is indeed radiative because the cooling time is much less than the age of the SN at both epochs e.g., days at days and days at days. Using these parameters and the equation above, we find erg s-1 at days and erg s-1 at days. Overall, both shock luminosities are consistent with the emergent shock powered emission observed in SN 2023ixf at both epochs and indicates nearly complete thermalization of RS X-rays converted to UVOIR emission by the CDS and inner ejecta. Intriguingly, the power budget in SN 2023ixf at days implies that the luminous near- and mid-IR emission from silicate and carbonaceous dust grains (e.g., Figures 1 & 8) is a product of shock powered emission since radioactive decay power is insufficient to heat the dust formed in the CDS and inner ejecta.
@@ 3.3 Radiative Transfer Model of the Low-Energy Radiation
374 | ||||||
620 |
As shown in Figure 7, the inclusion of dust can reproduce the asymmetries observed in both the boxy and narrow components of H, indicative of dust formation in the inner ejecta during the first complete spectral epoch ( day) as well as in the CDS by the next epoch ( days). Furthermore, the asymmetries in the line profiles appear to be wavelength dependent (e.g., see Fig. 6), which can be used to trace the dust opacity as has been done for type IIn SNe (e.g., SN 2010jl; Gall et al. 2014). Pre-existing dust external to the FS is likely excluded because it would cause significant attenuation at UV wavelengths, which is not observed in the HST spectra. This overall evolution of dust formation is supported by the detection of molecular CO emission beginning at days (Park et al., 2025; Medler et al., 2025). As shown in Appendix Figure A.2, we fit both epochs of IR spectra with analytic formalisms for dust emission (e.g., see Fox et al. 2010, 2011; Tinyanont et al. 2019; Shahbandeh et al. 2023; Pearson et al. 2025; Tinyanont et al. 2025)777https://github.com/stinyanont/sed_et_al after masking out all emission lines and find consistent silicate dust masses to what is adopted in the best-matched CMFGEN spectra. However, we find that the IR spectrum of SN 2023ixf requires an additional C-rich dust components to match the NIR emission, which is not included in the best-matched CMFGEN model spectra with Si-rich dust. Future simulations will explore combinations of both dust compositions.
In Figure 8, we present the best-matched CMFGEN models at days compared to the UVOIR SED of SN 2023ixf at days. We find that a shock power of erg s-1 is sufficient to match the total UV emission and the most prominent line profiles such as Ly and Mg ii as well as the boxy emission observed in H at days. At days, we find that the model shock power needs to be reduced to erg s-1 so as to not over-predict the UV emission lines, but also observe that the erg s-1 shock power model is a better match to the optical and NIR spectra at this phase. While the reduction in total UV flux between epochs is expected from a declining RS X-ray luminosity (e.g., Eqn. 1), we note that the observed shock powered emission is larger than the model prediction at days. Furthermore, at days, one CMFGEN model includes dust in the CDS and another with dust in the inner ejecta (, both of which have the most dramatic effect on the H and He emission line profiles. Further analysis including hybrid dust locations should be explored but is beyond the scope of this paper.

### v3

@@ Abstract
We present multi-wavelength observations of the type II supernova (SN II) 2023ixf during its first two years of evolution. We combine ground-based optical/NIR spectroscopy with Hubble Space Telescope (HST) far- and near-ultraviolet spectroscopy and James Webb Space Telescope (JWST) near- and mid-infrared photometry and spectroscopy to create spectral energy distributions of SN 2023ixf at +374 and +620 days post-explosion, covering a wavelength range of m. The multi-band light curve of SN 2023ixf follows a standard radioactive decay decline rate after the plateau until days, at which point shock powered emission from ongoing interaction between the SN ejecta and circumstellar material (CSM) begins to dominate. This evolution is temporally consistent with 0.3-10 keV X-ray detections of SN 2023ixf and broad “boxy” spectral line emission, which we interpret to signal reprocessing of shock luminosity in a cold dense shell located between forward and reverse shocks. Using the expected absorbed radioactive decay power and the detected X-ray luminosity, we quantify the total shock powered emission at the +374 and +620 day epochs and find that it can be explained by nearly complete thermalization of the reverse shock luminosity as SN 2023ixf interacts with a continuous, “wind-like” CSM with a progenitor mass-loss rate of M⊙ yr-1 (). Additionally, we construct multi-epoch spectral models from the non-LTE radiative transfer code CMFGEN, which contain radioactive decay and shock powers, as well as dust absorption, scattering, and emission. We find that models with shock powers of erg s-1 and of silicate dust in the cold dense shell and/or inner SN ejecta can effectively reproduce the global properties of the late-time ( days) UV-to-IR spectra of SN 2023ixf.
@@ 1 Introduction
Supernova (SN) 2023ixf is a type II supernova (SN II) that occurred on 2023-05-18 in Messier 101 at a distance of Mpc (Riess et al., 2022). SN 2023ixf is one of the closest SN in the last decade and, consequently, was reported to the Transient Name Server (TNS) within 1 day of first light by Koichi Itagaki (Itagaki, 2023) on 2023 May 19 17:27:15 (MJD 60083.73). The earliest reported detection of SN 2023ixf was on MJD 60082.85 at mag (-12.1 mag) in -band following a deep upper limit of mag ( mag) on MJD 60082.66 (Mao et al., 2023; Li et al., 2024). Consequently, an estimate of the time of first light is MJD , which is adopted throughout this work.
The early-time spectra of SN2023ixf exhibit emission lines from H i, He i/ii, C iv, and N iii/iv/v, superimposed on a hot, blue spectral continuum (Jacobson-Galán et al., 2023; Bostroem et al., 2023; Teja et al., 2023; Zhang et al., 2023). These lines, commonly referred to as “flash” or “IIn-like” features, likely originate from the sustained photoionization of dense circumstellar material (CSM) located ahead of the forward shock (Gal-Yam et al., 2014; Khazov et al., 2016; Yaron et al., 2017; Dessart et al., 2017; Terreran et al., 2022; Bruch et al., 2021, 2023; Dessart & Jacobson-Galán, 2023; Jacobson-Galán et al., 2024a). Much like Type IIn supernovae, the emission line profiles consist of a narrow core accompanied by broad Lorentzian “wings” that extend up to 1000 , a signature of electron scattering within the ionized, optically thick CSM (Chugai, 2001; Dessart et al., 2009, 2017; Huang & Chevalier, 2018). Additionally, SN 2023ixf has the earliest and highest cadence UV spectroscopy (Teja et al., 2023; Zimmerman et al., 2024), spectropolarimetry (Vasylyev et al., 2023; Singh et al., 2024; Shrestha et al., 2025; Vasylyev et al., 2025) and high resolution spectroscopy (Smith et al., 2023; Zimmerman et al., 2024; Dickinson et al., 2025) of any CSM-interacting SN II to date.
@@ 2.1 Photometric Observations
In addition to our observations, we include -band photometry from the Zwicky Transient Facility (ZTF; Bellm et al. 2019; Graham et al. 2019) forced-photometry service (Masci et al., 2019). We also include multi-band photometry presented in Jacobson-Galán et al. (2023); Van Dyk et al. (2024a); Hsu et al. (2024); Zimmerman et al. (2024); Singh et al. (2024) in our analysis. Furthermore, we include JWST MIRI imaging of SN 2023ixf from 2024-03-15 (JWST-GO-3921, PI: Fox) and 2025-01-27 (JWST-GO-5290, PI: Ashall), which includes observations in F770W, F1000W, F1500W, F1800W, F2100W, and F2500W filters. We performed PSF photometry on all filters and epochs using the spacephot photometry package and the PSF models from WebbPSF.333https://space-phot.readthedocs.io/en/latest/ All photometric measurements are consistent with those derived in Medler et al. (2025). The complete UVOIR light curve of SN 2023ixf is shown in Figure 2.
@@ 2.2 Spectroscopic Observations
Far- and near-UV spectra of SN 2023ixf were obtained on 2024-03-24 ( days) and 2025-01-26 ( days) with the Hubble Space Telescope (HST) using the Multi-Anode MicroChannel Array (MAMA) detectors on the Space Telescope Imaging Spectrograph (STIS) through HST programs 17497 (PI Valenti) and 17772 (PI Bostroem) Bostroem et al. (2024a, 2025). Standard bias subtraction, flat-fielding and cosmic ray rejection were performed automatically before downloading from the Mikulski Archive for Space Telescopes (MAST). Wavelength calibration, flux calibration and 1-D extraction was performed using the calstis pipeline routine. Near- and mid-infrared spectroscopy of SN 2023ixf were obtained on 2024-01-26, 2024-05-26, and 2025-01-08 ( days) (JWST-DD-4575, JWST-GO-5290; PI: Ashall) with the James Webb Space Telescope (JWST) using NIRSpec with G235M-F170LP and G395M-F290LP grating/filter combinations (Jakobsen et al., 2022) and MIRI/LRS (Kendrew et al., 2015). These data were reduced using the jwst555https://github.com/spacetelescope/jwst pipeline, which performs standard bias subtraction, flat-fielding, wavelength and flux calibrations, and spectral extraction. The complete optical/NIR spectral sequence is shown in Figure 3 and a log of all spectral observations utilized in this paper is provided in Table A1. All photometric and spectroscopic data will be made public online.666https://github.com/wynnjacobson-galan/SN2023ixf
@@ 3.1 Radioactive Decay Power
As shown in Figure 2, the photospheric phase persists as a linearly-declining plateau in the SN 2023ixf light curve for 83 days. Following the fall from the plateau, the photometric evolution of SN 2023ixf becomes consistent with radioactive decay from the standard Fe chain (Arnett, 1982). Modeling of the nebular decline rate has suggested and incomplete -ray trapping with a characteristic timescale of days (Zimmerman et al., 2024; Singh et al., 2024; Li et al., 2025). To confirm these parameters, we construct two variations of the post-plateau bolometric light curve, which are then fit with analytic formalisms for radioactive decay power. First, we construct a pseudobolometric light curve using filters and fit this light curve with the pseudobolometric light curve of SN 1987A derived with the same filter combination (e.g., see Eqn. 1 of Jacobson-Galán et al. 2025), which yields and days. We note that consistent values are found when this method is applied to the more confined filter combination. Second, we construct a pseudobolometric light curve from the combination of filters and which covers m. Given the emergence of shock powered emission at late-times (§3.2), we model this pseudobolometric light curve at days and find and days. We chose to exclude UV and mid-IR photometry in this light curve given that (1) of SN II luminosity is expected to be emitted between m at days (Dessart, 2025a) if it arises from decay power only and (2) the UV excess from persistent shock powered emission shown in Figure 5 would lead to an overestimate of . Based on the average of both modeling attempts, we choose to adopt and days when calculating the contribution of absorbed radioactive decay power to the observed luminosity of SN 2023ixf.
@@ 3.2 Shock Powered Emission
As shown in Figures 2 & 5, an excess emission beyond what is expected from pure radioactive decay first becomes apparent at UV wavelengths as the Swift-UVOT light curve becomes flat at days and remains bright at mag above the host galaxy background – these observations are also consistent with observed emission in the near- and far-UV at days (Bostroem et al., 2025). Host galaxy subtracted optical and NIR photometry shows a similar flattening at days and the pseudobolometric light curve (e.g., Fig. 5) deviates from a radioactive decay decline rate as shock powered emission from persistent SN ejecta-CSM interaction becomes the dominant power source in SN 2023ixf. Here we define “shock powered emission” as all additional UVOIR luminosity that does not come from absorbed radioactive decay power e.g., the reprocessed X-ray luminosity from the radiative RS. The presence of shock powered emission in SN 2023ixf is also confirmed by the spectral evolution shown in Figure 3 as the forbidden emission lines observed at year fade and prominent boxy H, H, P, He i and [Fe ii] emission become dominant at days – this transition in H is shown in Figure 6. The presence of these emission line profiles is direct evidence for a radiative reverse shock that injects power into both the CDS and the inner ejecta (Nymark et al., 2006; Chevalier & Fransson, 2017; Dessart & Hillier, 2022). Similarly, prominent Ly and Mg ii emission is observed at UV wavelengths, likely from the CDS (e.g., Fransson et al. 2005; Dessart et al. 2023; Bostroem et al. 2024a, 2025. Furthermore, evidence for ongoing CSM interaction is given by persistent X-ray and radio emission, both of which imply a mass loss rate of M⊙ yr-1 at shock radii cm (Panjkov et al., 2024; Chandra et al., 2024; Nayana et al., 2025). However, this shock powered emission source from ejecta interaction with a likely steady-state wind has remained subdominant compared to radioactive decay power until days in optical/IR bands and its emergence at these late times is not the result of SN ejecta interaction with some detached shell of CSM.
@@
where is the ejecta velocity at the RS in units of and is the CSM density power-law index. For a radiative cooling RS and a density profile, the time variable above disappears and the luminosity decreases as , which then goes as for and given the time dependence of the ejecta velocity (e.g., see Eqn. 2.2 in Fransson et al. 1996). We adopt M⊙ yr-1() and choose an ejecta velocity of , which covers the maximum velocity observed in the “boxy” H profile (e.g., Fig. 6). Applying Eqn. 3.14 in Fransson et al. (1996), we confirm that the RS is indeed radiative because the cooling time is much less than the age of the SN at both epochs e.g., days at days and days at days. Using these parameters and the equation above, we find erg s-1 at days and erg s-1 at days. At these phases, the shock powered emission () observed in SN 2023ixf is erg s-1 and erg s-1, both of which are consistent with the predicted RS emission. Overall, the emergent shock powered emission observed in SN 2023ixf indicates nearly complete thermalization of RS X-rays converted to UVOIR emission by the CDS and inner ejecta. Intriguingly, the power budget in SN 2023ixf at days implies that the luminous near- and mid-IR emission from silicate and carbonaceous dust grains (e.g., Figures 1 & 8) is a product of shock powered emission since radioactive decay power is insufficient to heat the dust formed in the CDS and inner ejecta.
@@ 3.3 Radiative Transfer Model of the Low-Energy Radiation
| 374 | ||||||
| 620 |
As shown in Figures 7 and A.3, the inclusion of dust can also produce similar blue/red asymmetries that are observed in both the boxy and narrow components of H, indicative of dust formation in the inner ejecta during the first complete spectral epoch ( day) as well as in the CDS by the next epoch ( days). Furthermore, the asymmetries in the line profiles may also be wavelength dependent (e.g., see Fig. 6), which can be used to trace the dust opacity as has been done for type IIn SNe (e.g., SN 2010jl; Gall et al. 2014) and could also suggest larger grain sizes of m. While some pre-existing dust external to the FS may exist, the overall CSM dust mass must be quite small because it would cause significant attenuation at UV wavelengths, which is not observed in the HST spectra. This overall evolution of dust formation is supported by the detection of molecular CO emission beginning at days (Park et al., 2025; DerKacy et al., 2025) and mid-infrared silicate emission (Medler et al., 2025). As shown in Appendix Figure A.2, we fit both epochs of IR spectra with analytic formalisms for dust emission (e.g., see Fox et al. 2010, 2011; Tinyanont et al. 2019; Shahbandeh et al. 2023; Pearson et al. 2025; Tinyanont et al. 2025)777https://github.com/stinyanont/sed_et_al after masking out all emission lines and find consistent silicate dust masses to what is adopted in the best-matched CMFGEN spectra. However, we find that the IR spectrum of SN 2023ixf requires an additional C-rich dust components to match the NIR emission, which is not included in the best-matched CMFGEN model spectra with Si-rich dust. Furthermore, it is clear from the analytic model fits that the emission at m is not well-matched – this could be due to a number of effects e.g., dust opacities, clumping, etc. (Sarangi, 2022; Dessart et al., 2025). Future simulations will explore combinations of various dust compositions, temperatures and structures.
In Figure 8, we present the best-matched CMFGEN models at days compared to the UVOIR SED of SN 2023ixf at days. We find that a shock power of erg s-1 is sufficient to match the total UV emission and the most prominent line profiles such as Ly and Mg ii as well as the boxy emission observed in H at days. At days, we find that the model shock power needs to be reduced to erg s-1 so as to not over-predict the UV emission lines, but also observe that the erg s-1 shock power model is a better match to the optical and NIR spectra at this phase. While the reduction in total UV flux between epochs is expected from a declining RS X-ray luminosity (e.g., Eqn. 1), we note that the observed shock powered emission is larger than the model prediction at days. Furthermore, at days, one CMFGEN model includes dust in the CDS and another with dust in the inner ejecta (, both of which have the most dramatic effect on the H and He emission line profiles. However, these current CMFGEN models show inconsistency with the spectral shape and overall flux in the IR, which is likely the product of a fixed dust temperature and single dust emission location within the SN. Further analysis including hybrid dust locations and dust temperature profiles should be explored but is beyond the scope of this paper.